A suspension of calcium carbonate particles in water flows through a pipe. Your assignment is to determine both the flow rate and the composition of this slurry. You proceed to collect the stream in a graduated cylinder for 1.00 min; you then weigh the cylinder, evaporate the collected water, and reweigh the cylinder. The following results are obtained:
Mass of empty cylinder: 65.0g
Mass of cylinder + collected slurry: 565g
Volume collected: 455mL
Mass of cylinder after evaporation: 215g
Calculate
(a) The volumetric flow rate and mass flow rate of the suspension.
(b) The density of the suspension.
(c) The mass fraction of CaCO3 in the suspension.
1
Expert's answer
2012-12-25T10:00:59-0500
(a) The volumetric flow rate
= 455 ml / min
The mass flow rate of the suspension
= (565-65= 500) = 500 g / min

(b) The density of the suspension
= 500 g / 455 ml
= 1.0989 g/ml ( 1.1)

(c) The mass fraction of CaCO3 in the suspension (mass of water is
the same that its volume 455 ml = 455 g )

w = (500-455) / 500 * 100% = 9 %